Can someone help me understand this please?
Solve each system by substitution. State the x-value of the solution.
x+4y=11
2x-5y=-4
A. -3 B. -2 C. 2 D. 3

Answer by ankor@dixie-net.com(22740)

Solve each system by substitution. State the x-value of the solution.
x + 4y = 11
2x- 5y = -4
:
We will use the 1st equation for substitution in the 2nd equation
x + 4y = 11
subtract 4y from both sides
x = -4y+11
:
2x - 5y = -4
substitute (-4y+11) for x in the 2nd equation
2(-4y+11) - 5y = -4
-8y + 22 - 5y = -4
-8y - 5y = -4 - 22
-13y = -26
y has to be positive, multiply both sides by -1
13y = 26
y = 26/13
y = 2
:
They want the value of x; use the equation: x - -4y+ 11, replace y with 2
x = -4(2) + 11
x = -8 + 11
x = 3
:
:
You can check these solutions in the original equations
Are you understanding this better now?